# The variational principle for a marked Gibbs point process with infinite-range multibody interactions

### v1

@@ 7. The support of good measures: Temperedness
$\Omega^{*}_{K}$ :=$:=\left\{\omega\in\Omega\ \middle|\ \forall n\geq K\ \forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}\leq g(n)\right\},$
$\Omega^{*}_{N,M}$ :=$:=\left\{\omega\in\Omega\ \middle|\ \forall n\in\mathbb{N}\cap[N,M]\ \forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}\leq g(n)\right\}.$
@@
$\pi^{z}((\Omega^{*}_{N,M})^{c})\lesssim_{\mathcal{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}$
@@
$\pi^{z}((\Omega^{*}_{K})^{c})=\sup_{M\geq K}\pi^{z}((\Omega^{*}_{K,M})^{c})\lesssim_{\mathcal{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}.$
@@
$=\exp\Big{(}-z\left\lvert\Lambda_{N}\right\rvert\mathcal{R}(R>g(N))-z\sum_{n=N+1}^{M}\left\lvert\Lambda_{n}\setminus\Lambda_{n-1}\right\rvert\mathcal{R}(R>g(n))\Big{)},$
@@
$\pi^{z}((\Omega^{*}_{N,M})^{c})$ $=1-\exp\Big{(}-z\left\lvert\Lambda_{N}\right\rvert\mathcal{R}(R>g(N))-z\sum_{n=N+1}^{M}\left\lvert\Lambda_{n}\setminus\Lambda_{n-1}\right\rvert\mathcal{R}\big{(}R>g(n)\big{)}\Big{)}$
$\leq 1-\exp\Big{(}-z\left\lvert\Lambda_{N}\right\rvert\mathcal{R}(R>g(N))-z\sum_{n=N+1}^{\infty}\left\lvert\Lambda_{n}\setminus\Lambda_{n-1}\right\rvert\mathcal{R}\big{(}R>g(n)\big{)}\Big{)}$
$\leq 1-\exp\Big{(}-c^{\prime}_{\mathcal{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}-c^{\prime\prime}_{\mathcal{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}\Big{)}\lesssim_{\mathcal{R},\gamma,z}\exp(-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}),$
@@
$\leq\frac{\log\big{(}c_{\mathcal{R},\gamma,z}\mathrm{e}^{\left\lvert\Lambda_{M}\right\rvert^{1+\gamma/2}}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}+1\big{)}}{\left\lvert\Lambda_{M}\right\rvert^{1+\gamma/2}}+\frac{I^{z}(P)}{\left\lvert\Lambda_{M}\right\rvert^{\gamma/2}}.$
@@
$P((\Omega^{*}_{N_{k},N_{k+1}})^{c})\lesssim\frac{\log\Big{(}c_{\mathcal{R},\gamma,z}\mathrm{e}^{-c_{d}k^{4(1+1/\gamma)}(1-k^{-2}((k+1)/k)^{4(1/2+1/\gamma)})}+1\big{)}}{k^{(4+1/2)d}}+\frac{I^{z}(P)}{k^{2d}}\lesssim k^{-2d}.$
@@
In order to show that infinite-volume Gibbs measures are tempered, we first provide a finite-volume comparison result.
@@
Since $\omega_{\Lambda}\mapsto\exp(-\beta H_{\Lambda,\zeta}(\omega_{\Lambda}))$ and $\omega_{\Lambda}\mapsto\mathbbm{1}\{\omega_{\Lambda}\in(\Omega_{\Lambda})^{*}_{K}\}$ are both decreasing, the result follows by applying the Harris-/FKG-inequality, i.e.,
@@
This finite-volume comparison now allows us to conclude that the infinite-volume Gibbs measures are also tempered.
@@
$\geq\int P(\mathrm{d}\zeta)G_{\Lambda,z,\beta,\zeta}\big{(}(\Omega_{\Lambda})_{K}^{*}\big{)}\geq\int P(\mathrm{d}\zeta)\pi^{z}(\Omega^{*}_{K})=\pi^{z}(\Omega^{*}_{K}).$
@@
[…]
@@
[…]
@@ 8.1. Periodic boundary conditions
at least if the periodized configuration $\omega^{(n)}$ respects the hardcore configuration. But otherwise both left- and right-hand side are $\infty$. ∎
@@
follows directly from [GZ93, Theorem 3.1]. The assumptions of the cited theorem are met since $H$ is trivially stable by being positive and lower-semicontinuous by Proposition 6.5 and we have $\big{|}\left\lvert\Lambda_{n}\right\rvert^{-1}\bigl{(}\beta H_{\Lambda_{n},\mathrm{per}}(\omega)-N_{\Lambda_{n}}\bigr{)}-\bigl{(}\beta H(R_{n,\omega})-R_{n,\omega}[N_{C}]\bigr{)}\big{|}=0$ by Lemma 8.1.
@@
does not directly follow from the mentioned theorem of Georgii–Zessin, but by inspection of its proof together with the fact that $H$ is continuous on $\{H<\infty\}$ (see Proposition 6.5) and the ergodic approximation given in Lemma 8.2 below. ∎
@@
$=P\left[\left(N_{x+\Delta}^{\psi}\right)^{1+\frac{d}{\delta}}\mathbbm{1}_{N_{x+\Delta}>1}\right]+P\left[\left(N_{x+\Delta}^{\psi}\right)^{1+\frac{d}{\delta}}\mathbbm{1}_{N_{x+\Delta}>1}\right]$
@@
Finally, as $H$ is continuous on the set $\{H<\infty\}$, and $H(P_{n})<\infty$, we have that $\limsup_{n\to\infty}I^{z}(P_{n})+H(P_{n})\leq I^{z}(P)+H(P)$. ∎
@@ 8.2. Free boundary conditions
$\pi^{z}\bigl{(}\exists\boldsymbol{x}\in\omega_{\Lambda_{K}}\colon R_{x}\geq g(K)\bigr{)}=\pi^{z}((\Omega^{*}_{K,K})^{c})\lesssim\mathrm{e}^{-\left\lvert\Lambda_{K}\right\rvert^{1+\gamma}}$
by Lemma 7.1, which we will use in the following to establish limits.
Consider the events $A_{n}=\{\omega\in\Omega_{\Lambda_{n}}\,|\,\omega_{\Lambda_{n}\setminus\Lambda_{n-g(n)}}=\emptyset\}$ and $B_{n}=\{\omega\in\Omega_{\Lambda_{n}}\,|\,\forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}<g(n)\}$. Then, on $A_{n}\cap B_{n}$, we have $H_{\Lambda_{n},\mathrm{per}}=H_{\Lambda_{n},\emptyset}=H_{\Lambda_{n-g(n)},\emptyset}$ and therefore
$\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\Big{]}$ $\geq\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{B_{n}}\Big{]}$
$=\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n-g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{B_{n}}\Big{]}$
$=\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n-g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\Big{]}-\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n-g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{B_{n}^{c}}\Big{]}$
$=\mathrm{e}^{z\left\lvert\Lambda_{n}\right\rvert-z\left\lvert\Lambda_{n-g(n)}\right\rvert}\pi^{z}_{\Lambda_{n-g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n-g(n)},\emptyset}}\Big{]}-\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n-g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{B_{n}^{c}}\Big{]}.$
It follows that
$\liminf_{n\to\infty}\frac{\log(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\right])}{|\Lambda_{n}|}$ $\geq\liminf_{n\to\infty}\frac{\log\Big{(}\pi^{z}_{\Lambda_{n-g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n-g(n)},\emptyset}}\Big{]}\Big{)}}{|\Lambda_{n-g(n)}|}$
$\geq\liminf_{n\to\infty}\frac{\log(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\right])}{|\Lambda_{n}|}.$
Now let $A^{\prime}_{n}=\{\omega\in\Omega_{\Lambda_{n+g(n)}}\,|\,\omega_{\Lambda_{n+g(n)}\setminus\Lambda_{n}}=\emptyset\}$ and $B^{\prime}_{n}=\{\omega\in\Omega_{\Lambda_{n+g(n)}}\,|\,\forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}<g(n)\}$. Then, we have
$\pi^{z}_{\Lambda_{n+g(n)}}$ $\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\mathrm{per}}}\Big{]}\geq\pi^{z}_{\Lambda_{n+g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\mathrm{per}}}\mathbbm{1}_{A^{\prime}_{n}}\mathbbm{1}_{B^{\prime}_{n}}\Big{]}$
$=\pi^{z}_{\Lambda_{n+g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A^{\prime}_{n}}\mathbbm{1}_{B^{\prime}_{n}}\Big{]}$
$=\pi^{z}_{\Lambda_{n+g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A^{\prime}_{n}}\Big{]}-\pi^{z}_{\Lambda_{n+g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A^{\prime}_{n}}\mathbbm{1}_{{B^{\prime}_{n}}^{c}}\Big{]}$
$=\mathrm{e}^{z\left\lvert\Lambda_{n+g(n)}\right\rvert-z\left\lvert\Lambda_{n}\right\rvert}\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\Big{]}-\pi^{z}_{\Lambda_{n+g(n)}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A^{\prime}_{n}}\mathbbm{1}_{{B^{\prime}_{n}}^{c}}\Big{]}.$
It follows that
@@
Finally,
$\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\mathbbm{1}_{B_{n}}$ $\geq\Big{(}\mathrm{e}^{\beta c\big{(}\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n-g(n)}\right\rvert\big{)}}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{H_{\Lambda_{n},\mathrm{per}}<\infty}+0\cdot\mathbbm{1}_{H_{\Lambda_{n},\mathrm{per}}=\infty}\Big{)}\mathbbm{1}_{B_{n}}$
$=\mathrm{e}^{\beta c\big{(}\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n-g(n)}\right\rvert\big{)}}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{B_{n}}$
@@
$\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\Big{]}\geq\mathrm{e}^{\beta c\big{(}\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n-g(n)}\right\rvert\big{)}}$ $\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\right]+\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\mathbbm{1}_{B_{n}^{c}}\Big{]}$
$-\mathrm{e}^{\beta c\big{(}\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n}-g(n)\right\rvert\big{)}}\pi^{z}_{\Lambda_{n}}\Big{[}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{B_{n}^{c}}\Big{]},$
and thus
$\liminf_{n\to\infty}\frac{\log\big{(}\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\right]\big{)}}{|\Lambda_{n}|}\geq\liminf_{n\to\infty}\frac{\log\big{(}\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\right]\big{)}}{|\Lambda_{n}|},$
and also
$\limsup_{n\to\infty}\frac{\log\big{(}\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\right]\big{)}}{|\Lambda_{n}|}\geq\limsup_{n\to\infty}\frac{\log\big{(}\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\right]\big{)}}{|\Lambda_{n}|},$

### v2

@@ 7. The support of good measures: Temperedness
$\Omega^{*}_{K}$ $:=\left\{\omega\in\Omega\ \middle|\ \forall n\geq K\ \forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}\leq g(n)\right\},$
$\Omega^{*}_{N,M}$ $:=\left\{\omega\in\Omega\ \middle|\ \forall n\in\mathbb{N}\cap[N,M]\ \forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}\leq g(n)\right\}.$
@@
$\pi^{z}((\Omega^{*}_{N,M})^{c})\lesssim_{\mathfrak{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}$
@@
$\pi^{z}((\Omega^{*}_{K})^{c})=\sup_{M\geq K}\pi^{z}((\Omega^{*}_{K,M})^{c})\lesssim_{\mathfrak{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}.$
@@
$=\exp\Big(-z\left\lvert\Lambda_{N}\right\rvert\mathfrak{R}(R>g(N))-z\sum_{n=N+1}^{M}\left\lvert\Lambda_{n}\setminus\Lambda_{n-1}\right\rvert\mathfrak{R}(R>g(n))\Big),$
@@
$\pi^{z}((\Omega^{*}_{N,M})^{c})$ $=1-\exp\Big(-z\left\lvert\Lambda_{N}\right\rvert\mathfrak{R}(R>g(N))-z\sum_{n=N+1}^{M}\left\lvert\Lambda_{n}\setminus\Lambda_{n-1}\right\rvert\mathfrak{R}\big(R>g(n)\big)\Big)$
$\leq 1-\exp\Big(-z\left\lvert\Lambda_{N}\right\rvert\mathfrak{R}(R>g(N))-z\sum_{n=N+1}^{\infty}\left\lvert\Lambda_{n}\setminus\Lambda_{n-1}\right\rvert\mathfrak{R}\big(R>g(n)\big)\Big)$
$\leq 1-\exp\Big(-c^{\prime}_{\mathfrak{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}-c^{\prime\prime}_{\mathfrak{R},\gamma,z}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}\Big)\lesssim_{\mathfrak{R},\gamma,z}\exp(-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}),$
@@
$\leq\frac{\log\big(c_{\mathfrak{R},\gamma,z}\mathrm{e}^{\left\lvert\Lambda_{M}\right\rvert^{1+\gamma/2}}\mathrm{e}^{-\left\lvert\Lambda_{N}\right\rvert^{1+\gamma}}+1\big)}{\left\lvert\Lambda_{M}\right\rvert^{1+\gamma/2}}+\frac{I^{z}(P)}{\left\lvert\Lambda_{M}\right\rvert^{\gamma/2}}.$
@@
$P((\Omega^{*}_{N_{k},N_{k+1}})^{c})\lesssim\frac{\log\Big(c_{\mathfrak{R},\gamma,z}\mathrm{e}^{-c_{d}k^{4(1+1/\gamma)}(1-k^{-2}((k+1)/k)^{4(1/2+1/\gamma)})}+1\big)}{k^{(4+1/2)d}}+\frac{I^{z}(P)}{k^{2d}}\lesssim k^{-2d}.$
@@
In order to show that infinite-volume translation-invariant Gibbs measures are tempered, we first provide a finite-volume comparison result.
@@
Since $\omega_{\Lambda}\mapsto\exp(-\beta H_{\Lambda,\zeta}(\omega_{\Lambda}))$ and $\omega_{\Lambda}\mapsto\mathbbm{1}\{\omega_{\Lambda}\in(\Omega_{\Lambda})^{*}_{K}\}$ are both decreasing, the result follows by applying the Harris-/FKG-inequality for the Poisson point process $\pi^{z}_{\Lambda}$, i.e.,
@@
This finite-volume comparison now allows us to conclude that the translation-invariant infinite-volume Gibbs measures are also tempered.
@@
$\geq\int P(\mathrm{d}\zeta)\pi^{z}(\Omega^{*}_{K})=\pi^{z}(\Omega^{*}_{K}).$
@@
$\nu_{\Lambda}(A)=\int P(\mathrm{d}\zeta)\int G_{\Lambda,z,\beta,\zeta}(\mathrm{d}\omega_{\Lambda})\,\mathbbm{1}\{\omega_{\Lambda}\zeta_{\Lambda^{c}}\in A\}$
@@
$=\int\,P(\mathrm{d}\zeta)\,\mathbbm{1}\{\zeta\in\Omega^{*}_{K}\}\,G_{\Lambda,z,\beta,\zeta}\big((\Omega_{\Lambda})^{*}_{K}\big)$
$\geq\int\,P(\mathrm{d}\zeta)\,\mathbbm{1}\{\zeta\in\Omega^{*}_{K}\}\,\pi^{z}(\Omega^{*}_{K})$
@@ 8.1. Periodic boundary conditions
at least if the periodized configuration $\omega^{(n)}$ respects the hardcore condition. But otherwise both left- and right-hand side are $\infty$. ∎
@@
follows directly from [GZ93, Theorem 3.1]. The assumptions of the cited theorem are met since $H$ is trivially stable by being positive and lower-semicontinuous by Proposition 6.5 and we have $\big|\left\lvert\Lambda_{n}\right\rvert^{-1}\bigl(\beta H_{\Lambda_{n},\mathrm{per}}(\omega)-\log(z)N_{\Lambda_{n}}\bigr)-\bigl(\beta H(R_{n,\omega})-\log(z)R_{n,\omega}[N_{C}]\bigr)\big|=0$ by Lemma 8.1.
@@
does not directly follow from the mentioned theorem of Georgii–Zessin, but it follows from a closer inspection of its proof, together with the fact that $H$ is continuous on $\{H<\infty\}$ (see Proposition 6.5) and the ergodic approximation given in Lemma 8.2 below. ∎
@@
$=P\left[\left(N_{x+\Delta}^{\psi}\right)^{1+\frac{d}{\delta}}\mathbbm{1}_{N_{x+\Delta}\leq 1}\right]+P\left[\left(N_{x+\Delta}^{\psi}\right)^{1+\frac{d}{\delta}}\mathbbm{1}_{N_{x+\Delta}>1}\right]$
@@
Finally, as $H$ is continuous on the set $\{H<\infty\}$, and $H(P^{(n)})<\infty$, we have that $\limsup_{n\to\infty}I^{z}(P^{(n)})+H(P^{(n)})\leq I^{z}(P)+H(P)$. ∎
@@ 8.2. Free boundary conditions
$\pi^{z}\bigl(\exists\boldsymbol{x}\in\omega_{\Lambda_{K}}\colon R_{x}\geq g(K)\bigr)=\pi^{z}((\Omega^{*}_{K,K})^{c})\lesssim\mathrm{e}^{-\left\lvert\Lambda_{K}\right\rvert^{1+\gamma}}$ (8.1)
by Lemma 7.1, which we will use in the following to establish limits. Consider the events $A_{n}=\{\omega\in\Omega_{\Lambda_{n+g(n)}}\,|\,\omega_{\Lambda_{n+g(n)}\setminus\Lambda_{n}}=\emptyset\}$ and $B_{n}=\{\omega\in\Omega_{\Lambda_{n+g(n)}}\,|\,\forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}<g(n)\}$. Then, on $A_{n}\cap B_{n}$, we have $H_{\Lambda_{n+g(n)},\mathrm{per}}=H_{\Lambda_{n+g(n)},\emptyset}=H_{\Lambda_{n},\emptyset}$ and therefore
$\pi^{z}_{\Lambda_{n+g(n)}}$ $\Big[\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\mathrm{per}}}\Big]\geq\pi^{z}_{\Lambda_{n+g(n)}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\mathrm{per}}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{B_{n}}\Big]$
$=\pi^{z}_{\Lambda_{n+g(n)}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{B_{n}}\Big]$
$=\pi^{z}_{\Lambda_{n+g(n)}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\Big]-\pi^{z}_{\Lambda_{n+g(n)}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{{B_{n}}^{c}}\Big]$
$=\mathrm{e}^{z\left\lvert\Lambda_{n+g(n)}\right\rvert-z\left\lvert\Lambda_{n}\right\rvert}\pi^{z}_{\Lambda_{n}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\Big]-\pi^{z}_{\Lambda_{n+g(n)}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n+g(n)},\emptyset}}\mathbbm{1}_{A_{n}}\mathbbm{1}_{{B_{n}}^{c}}\Big]$
It follows, using the estimate 8.1, that
@@
Finally, considering the event $B^{\prime}_{n}=\{\omega\in\Omega_{\Lambda_{n}}\,|\,\forall\boldsymbol{x}\in\omega_{\Lambda_{n}}\colon R_{x}<g(n)\}$, we have
$\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\mathbbm{1}_{B^{\prime}_{n}}$ $\geq\Big(\mathrm{e}^{\beta c\big(\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n-g(n)}\right\rvert\big)}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{H_{\Lambda_{n},\mathrm{per}}<\infty}+0\cdot\mathbbm{1}_{H_{\Lambda_{n},\mathrm{per}}=\infty}\Big)\mathbbm{1}_{B^{\prime}_{n}}$
$=\mathrm{e}^{\beta c\big(\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n-g(n)}\right\rvert\big)}\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{B^{\prime}_{n}}$
@@
$\pi^{z}_{\Lambda_{n}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\Big]\geq\mathrm{e}^{\beta c\big(\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n-g(n)}\right\rvert\big)}$ $\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\right]+\pi^{z}_{\Lambda_{n}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\mathbbm{1}_{{B^{\prime}_{n}}^{c}}\Big]$
$-\mathrm{e}^{\beta c\big(\left\lvert\Lambda_{n}\right\rvert-\left\lvert\Lambda_{n}-g(n)\right\rvert\big)}\pi^{z}_{\Lambda_{n}}\Big[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\mathbbm{1}_{{B^{\prime}_{n}}^{c}}\Big],$
and thus, using again the estimate 8.1,
$\liminf_{n\to\infty}\frac{\log\big(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\emptyset}}\right]\big)}{|\Lambda_{n}|}\geq\liminf_{n\to\infty}\frac{\log\big(\pi^{z}_{\Lambda_{n}}\left[\mathrm{e}^{-\beta H_{\Lambda_{n},\mathrm{per}}}\right]\big)}{|\Lambda_{n}|}$